OpenStudy (anonymous):

A submarine is diving to a total depth of 45 feet. It descends 15% of the total depth in 1 minute. How deep will the submarine be after 4 minutes? −6.75 feet −27 feet −13.5 feet −20.25 feet

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (anonymous):

help ill give medal and fan

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (dangerousjesse):

What's \(\large 15\times 4\)?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (dangerousjesse):

So what's \(\large 45\times 1.60\)?

OpenStudy (anonymous):

72

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (dangerousjesse):

Ignore that -.-

OpenStudy (anonymous):

oh ok lol

OpenStudy (anonymous):

yea cus thats not one of the answer options

OpenStudy (anonymous):

XD

OpenStudy (dangerousjesse):

\(large 45 \times .15=x\) Solve for x :P

OpenStudy (dangerousjesse):

I meant \(\large 45 \times .15=x\)

OpenStudy (anonymous):

6.75

OpenStudy (anonymous):

thanks lol

OpenStudy (dangerousjesse):

And \(\large 6.75 \times 4\)?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

wait so the answer isent 6.75

OpenStudy (anonymous):

its 27

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (dangerousjesse):

Yes, it's 27 :) \(\large 6.75=\frac{1}{4}\) of the answer.